# Betelgeuse, Betelgeuse, Betelgeuse, Betel-buddy? Constraints on the dynamical companion to  Orionis from HST

### v1

@@ 2.2 Motivation to Observe at Quadrature
The luminosity ratio between Betelgeuse and in the YSO scenario is expected to be of order â, making undetectable in integrated light (MacLeod et al., 2025). However, in the far-UV spectrum of both YSOs and RSGs, emission from the chromosphere dominates the photosphere continuum (Linsky, 2017). HST has resolved many chromospheric lines in the spectra of YSOs (Yang et al., 2012; Roman-Duval et al., 2020, 2025), and has observed the chromosphere of Betelgeuse (Carpenter et al., 2018). This leads to three possible means of detection for : (1) the detection of a slight UV excess relative to observations taken where the companion is not present, (2) strong emission lines either shifted relative to, or not expected from, the chromosphere of Betelgeuse, and (3) RV variations in those UV lines over the timescale of an LSP period. Option (3) would require continuous observation over the 5.5-year duration of an LSP, which is impractical. As MacLeod et al. (2025) point out, option (1) is tenuous given intrinsic variations in Betelgeueseâs bright chromosphere. However, if observed at a proper orbital configuration where âs emission is redshifted or blueshifted relative to Betelgeuse, option (2) could be promising (whether or not the companion is coeval with Betelgeuse, and whether or not it is indeed a YSO).
We therefore target our observations such that âs emission is expected to be significantly redshifted relative to Betelgeuseâs, near receding quadrature on 6 December 2024 (Goldberg et al. 2024, within of the MacLeod et al. (2025) ephemeris which places quadrature slightly earlier in mid-to-late November 2024). We also note that the companionâs next (approaching) quadrature is years down the line (late November 2027) after passing behind the disc of Betelgeuse.
@@ 2.3 Far UV Expectations for a YSO Companion
In the far ultra-violet (FUV), the overall flux of Betelgeuse is expected to dominate over that of a YSO companion. Figure 1 shows the Color-Magnitude Diagram (CMD) for all Gaia stars with GALEX FUV photometry (â1,750âĂ ) across galactic latitudes â and longitudes â within âpc, which encompasses the 8â15âMyr old Scorpius-Centaurus (Sco-Cen) OB association. The official Sco-Cen members from Luhman (2022) are shown in blue. For both main sequence (MS) and 10âMyr pre-MS stars (excluding the white dwarf sequence to the bottom left), the Gaia (BPâRP) color roughly corresponds to stellar mass, indicated on the top x-axis. For (A/F stars), the field and Sco-Cen stars follow the same MS relation. G-type () and K-type () stars exhibit FUV excess compared to their field MS counterparts. Early M-stars () exhibit a wide range of FUV fluxes and chromospheric activity, some of them even brighter than the GK stars. This factor-of-ten spread in FUV brightness is expected, as emission lines and FUV flux depend on a starâs chromospheric activity, which varies with rotation and time and environmental properties in addition to mass and age.
@@
The weighted average FUV flux of Betelgeuse (135â175ânm) is âergâs-1âcm-2âĂ -1 (Carpenter et al., 2018), which corresponds to an apparent brightness of âmag and absolute brightness of âmag in the GALEX passband given the 168âpc distance to Betelgeuse (Joyce et al., 2020). Horizontal lines indicating 100%, 10%, and 1% of Betelgeuseâs FUV brightness are shown on the CMD (Fig. 1). Based on the available E140M spectrum of Betelgeuse from Carpenter et al. (2018), a companion with 30% the FUV flux of Betelgeuse should have already been detected.
@@
For possible masses of in the range, the companionâs luminosity in the FUV is expected to be of that of Betelgeuse, making it difficult to detect a shift in the overall flux level. However, the chromospheres of stars and YSOs in the Far UV are dominated by emission lines (see Linsky 2017 for a review). As discussed above, targeting our observations near the companionâs receding quadrature corresponds to Betelgeuseâs minimal RV (maximum Betelgeuse blueshift), but the companionâs maximal RV (maximum redshift). This orbital configuration entails that the emission features of should be redshifted with respect to the same features of Betelgeuse, by a speed of 45.5 km/s. This is well above the detectable limit for HST-STIS with a good resolution grating (10 km/s). Therefore, for emission lines that are sufficiently strong in both the chromospheres of Betelgeuse and YSOs, a second, redshifted emission line should be visible.
@@
For want of first-principles chromospheric modeling, reasonable proxies for the possible FUV spectra of are provided by the ULLYSES project, which has compiled, in a publicly available database, the HST UV spectroscopy of many pre-MS/T Tauri stars (Roman-Duval et al., 2020, 2025). To illustrate this, we choose five YSOs from the ULLYSES dataset333https://ullyses.stsci.edu/ullyses-targets-ttauri.html across the range of relevant masses and ages â CVSO 104 (0.77Mâ, 11 Myr), CVSO 165 (0.91Mâ, 11 Myr), MP Mus (1.1Mâ, 15 Myr), HD1353344B (1.5Mâ, 9 Myr), and HD104237 (2.0Mâ, 5 Myr). Later in this paper, we also use two other YSOs for comparative purposes â PDS 70 (0.76Mâ, 11 Myr), and AK Sco (1.52Mâ, 11Myr). The masses are obtained from the ULLYSES website and the ages from estimates of the age of the cluster containing the YSO (Blaauw, 1964; de Geus et al., 1989; Mamajek et al., 2002; MĂŒller et al., 2011; Bossini et al., 2019; Esposito et al., 2020). These objects additionally represent a wide range of .
@@
These data, along with the HST/STIS E140M spectrum of Betelgeuse from Carpenter et al. (2018), are shown in Figure 2. At the time of the Carpenter et al. (2018) observations (conducted in 2011), was transiting Betelgeuse, nearly at the middle of transit (see also § 3.4). Thus, while the overall flux of each emission line would contain contributions from both Betelgeuse and , their relative radial velocities should be close to zero â the only RV shift present at that time should be the overall peculiar motion of the system (22 km/s) as well as any RV changes induced by Betelgeuseâs pulsational or convective activity. In Figure 2, these data are shown as a thick solid orange line (labeled  Ori 2011), and have been blueshifted by km/s to match the expected total RV of Betelgeuse during our 2024 observations. We likewise shift the ULLYSES YSO spectra to be at the expected radial velocity of during the 2024 observations, 45.5 km/s redshifted. They have also been corrected for distance (168 pc, Joyce et al. 2020) and extinction ( and , MontargĂšs et al. 2021) appropriate to the parameters of the Betelgeuse system.
@@
While the overall flux in the UV of these YSOs is only 1â5% of that of Betelgeuse, several features could be visible in a STIS exposure containing (as opposed to exposures containing only the chromosphere of Betelgeuse). There are 8 strong emission lines in the ULLYSES YSO spectra that are close to or as strong as the same line from Betelgeuse, and 4 that do not appear as emission lines in the spectrum of Betelgeuse. The four panels of Figure 2 highlight four wavelength ranges with promising features that may be visible in our observations of the system.444See Figure 7 and Table 2 of Linsky (2017) for the full suite of lines we examined here and in the observing proposal.
@@
We also note that the ULLYSES data are all of isolated YSO systems. YSOs are often surrounded by, and accrete from, circumstellar disks, leading to the emission âwingsâ seen in some of the ULLYSES spectra. These are protoplanetary disks at early (10âMyr) times, and gas-poor processed disks at later (10âMyr) times (Ribas et al., 2015). These disks eventually photo-evaporate from high-energy radiation from the central star (Gorti & Hollenbach, 2009). However, these circumstellar disks can also be photoevaporated by UV radiation from nearby OB stars (Adams et al., 2004; Andrews, 2020), which has been directly observed for YSOs in young OB associations (Johnstone et al., 1998). For binary YSOs, systems with circumprimary but no circumsecondary disks are also much more common than the inverse (White & Ghez, 2001), and Kraus et al. (2012) finds that 2/3rds of YSO systems with close separations (r40âAU, compared to the 9AU separation for ) disperse their disks with 1âMyr of formation. Given that the  Oriâ system has a small separation relative to 40AU, and given that Betelguse was an extremely hot O star during its main-sequence lifetime, there is a high likelihood that a circumstellar disk around the companion has been dispersed by present day. Thus, the wide emission features stemming from strong winds or accretion (red/blue-shifted relative to the companion) that are common in YSO spectra around CII, and CIV lines (seen in the ULLYSES spectra in panels (a) and (d) of Figure 2), as well as wings around Lyman-, would almost certainly not be present in the  Ori system.
@@
We therefore focus our efforts on narrow lines without those wings, such as fluorescent H2 lines (seen in all four panels of Figure 2), as well as Lyman (discussed in §4.3). Of the spectra shown here, the 2âMâ YSO would be plainly visible in these lines, while the 0.77âMâ YSO would never be visible even with a large RV offset between Betelgeuse and . The 0.91âMâ YSO also appears generally too faint to detect. The 1.1 and 1.5âMâ YSOs are more promising to identify, though the extra flux added to the system is small. We also note the important caveat that these are individual, empirical YSO observations, not theoretical models of YSO FUV chromospheric features across the full range of possible companion masses, for which first-principles calculations remain elusive.
@@ 3 Observations with HST
Below, we summarize the observational history of Betelgeuse to place our campaign in context. We then describe the parameters of our own observational campaign that took place in November 2024. We emphasize the importance of the timing of our observations relative to the companionâs orbital phase (receding quadrature). We also describe our unique aperture tiling pattern, which allows for co-incident spectra which both do and do not contain emission from the companion.
@@ 3.1 Past UV Observations
Long before the launch of HST, early measures of the ultraviolet spectrum of Betelgeuse (spectral type M1-M2) were obtained from a balloon-borne telescope-spectrometer, the OAOâ2 spacecraft (Wisconsin photometer-spectrometer), and OAOâ3 (Princeton experiment on Copernicus). All detected the Mg II doublet near 2800âĂ , a strong emission doublet in the near-ultraviolet (Kondo et al., 1972; Doherty, 1972; Bernat & Lambert, 1976). Subsequently, the International Ultraviolet Explorer (IUE) extended spectroscopy further into the ultraviolet to the Lyman- line near 1200âĂ . These low resolution spectra show no evidence of any species of ionization beyond Si II and Fe II, suggesting electron temperatures less than about 10,000âK (Basri et al., 1981). The Hopkins Ultraviolet Telescope on the Space Shuttle offered the possibility of spectroscopy from 800âĂ Â to 1800âĂ Â (Davidsen, 1993); however, the few spectra of Betelgeuse were too low signal-to-noise to be conclusive.
@@
A substantial advance in the study of the near ultraviolet spectrum came with the Goddard High Resolution Spectrograph (GHRS) on HST, acquiring a detailed spectrum from 1980â3000âĂ Â that enabled the identification of many features (Brandt et al., 1995) from single and doubly ionized species, among them Fe II, C II, Ni II, Ti II in addition to molecular emission of OH. The GHRS instrument also accessed the far ultraviolet spectrum, from 1200âĂ Â to 1930âĂ , a region that is dominated by CO bandhead absorption in addition to the singly and doubly ionized species, C I, S I, O I, Cr II, and Si II (Carpenter et al., 1994). Those authors also claim the presence of a true continuum, with a color temperature of 3500â4800âK which they ascribe to the chromosphere, and suggest an origin in hot chromospheric regions. A large hot region was later observed on the extended ultraviolet disk in the 2500âĂ Â continuum, with temperatures in excess of 200âK from the disk value (Gilliland & Dupree, 1996), thus supporting origin in a hot region. Carpenter et al. (1994) in fact consider but dismiss the idea that a companion creates the continuum, in part because the observed flux distribution of warmer stars does not match the continuum shape.
@@
The spectral range below 1200âĂ , became generally accessible with the launch of the Far Ultraviolet Spectroscopic Explorer (FUSE), which enabled access from 905â1179âĂ Â where highly ionized species are expected. However, a spectrum of Betelgeuse revealed no sign of C III or O VI which are found in giants and supergiants of spectral type F, G, and K (Dupree et al., 2005).
@@
More recently, observations of Betelgeuse (PI T. Ayres 2010) were performed with STIS E140M using a square âarcsec aperture for 4.5 hours (Carpenter et al., 2018).The dates of these observations were February 7, 8, 24, 26, and April 1st 2011, at which point was transiting (passing in front of) Betelgeuse. While our campaign used the same instrument, grating, and aperture as the 2011 campaign, ours differs critically in timing, tiling, and motivation. The previous Ayres (2010); Carpenter et al. (2018) campaign was only interested in the overall UV spectrum of the chromosphere, and they did not vary their pointing. Therefore, that campaign did not obtain any spatial information and did not collect any âcleanâ spectra without .

### v2

@@ 2.2 Motivation to Observe at Quadrature
The luminosity ratio between Betelgeuse and in the YSO scenario is expected to be of order –, making undetectable in integrated light (MacLeod et al., 2025). However, in the far-UV spectrum of both YSOs and RSGs, emission from the chromosphere dominates the photosphere continuum (Linsky, 2017). HST has resolved many chromospheric lines in the spectra of YSOs (Yang et al., 2012; Roman-Duval et al., 2020, 2025), and has observed the chromosphere of Betelgeuse (Carpenter et al., 2018). This leads to three possible means of detection for : (1) the detection of a slight UV excess relative to observations taken where the companion is not present, (2) strong emission lines either shifted relative to, or not expected from, the chromosphere of Betelgeuse, and (3) RV variations in those UV lines over the timescale of an LSP period. Option (3) would require continuous observation over the 5.5-year duration of an LSP, which is impractical. As MacLeod et al. (2025) point out, option (1) is tenuous given intrinsic variations in Betelgeuese’s bright chromosphere. However, if observed at a proper orbital configuration where ’s emission is redshifted or blueshifted relative to Betelgeuse, option (2) could be promising (whether or not the companion is coeval with Betelgeuse, and whether or not it is indeed a YSO).
We therefore target our observations such that ’s emission is expected to be significantly redshifted relative to Betelgeuse’s, near receding quadrature on 6 December 2024 (Goldberg et al. 2024, within of the MacLeod et al. (2025) ephemeris which places quadrature slightly earlier in mid-to-late November 2024). We also note that the companion’s next (approaching) quadrature is years down the line (late November 2027) after passing behind the disc of Betelgeuse.
@@ 2.3 Far UV Expectations for a YSO Companion
In the far ultra-violet (FUV), the overall flux of Betelgeuse is expected to dominate over that of a YSO companion. Figure 1 shows the Color-Magnitude Diagram (CMD) for all Gaia stars with GALEX FUV photometry (–1,750 Å) across galactic latitudes – and longitudes – within pc, which encompasses the 8–15 Myr old Scorpius-Centaurus (Sco-Cen) OB association. The official Sco-Cen members from Luhman (2022) are shown in blue. For both main sequence (MS) and 10 Myr pre-MS stars (excluding the white dwarf sequence to the bottom left), the Gaia (BP–RP) color roughly corresponds to stellar mass, indicated on the top x-axis. For (A/F stars), the field and Sco-Cen stars follow the same MS relation. G-type () and K-type () stars exhibit FUV excess compared to their field MS counterparts. Early M-stars () exhibit a wide range of FUV fluxes and chromospheric activity, some of them even brighter than the GK stars. This factor-of-ten spread in FUV brightness is expected, as emission lines and FUV flux depend on a star’s chromospheric activity, which varies with rotation and time and environmental properties in addition to mass and age.
@@
The weighted average FUV flux of Betelgeuse (135–175 nm) is erg s-1 cm-2 Å-1 (Carpenter et al., 2018), which corresponds to an apparent brightness of mag and absolute brightness of mag in the GALEX passband given the 168 pc distance to Betelgeuse (Joyce et al., 2020). Horizontal lines indicating 100%, 10%, and 1% of Betelgeuse’s FUV brightness are shown on the CMD (Figure 1). Based on the available E140M spectrum of Betelgeuse from Carpenter et al. (2018), a companion with 30% the FUV flux of Betelgeuse should have already been detected.
@@
For possible masses of in the range, the companion’s luminosity in the FUV is expected to be of that of Betelgeuse, making it difficult to detect a shift in the overall flux level. However, the chromospheres of stars and YSOs in the Far UV are dominated by emission lines (see Linsky 2017 for a review). As discussed above, targeting our observations near the companion’s receding quadrature corresponds to Betelgeuse’s minimal RV (maximum Betelgeuse blueshift), but the companion’s maximal RV (maximum redshift). This orbital configuration entails that the emission features of should be redshifted with respect to the same features of Betelgeuse, by a speed of 45.5 km/s. This is well above the detectable limit for HST-STIS with a good resolution grating (10 km/s). Therefore, for emission lines that are sufficiently strong in both the chromospheres of Betelgeuse and YSOs, a second, redshifted emission line should be visible.
@@
For want of first-principles chromospheric modeling, reasonable proxies for the possible FUV spectra of are provided by the ULLYSES project, which has compiled, in a publicly available database, the HST UV spectroscopy of many pre-MS/T Tauri stars (Roman-Duval et al., 2020, 2025). To illustrate this, we choose five YSOs from the ULLYSES dataset333https://ullyses.stsci.edu/ullyses-targets-ttauri.html across the range of relevant masses and ages – CVSO 104 (0.77M⊙, 11 Myr), CVSO 165 (0.91M⊙, 11 Myr), MP Mus (1.1M⊙, 15 Myr), HD1353344B (1.5M⊙, 9 Myr), and HD104237 (2.0M⊙, 5 Myr). Later in this paper, we also use two other YSOs for comparative purposes – PDS 70 (0.76M⊙, 11 Myr), and AK Sco (1.52M⊙, 11Myr). The masses, distances, and values are obtained from the ULLYSES website and the ages from estimates of the age of the cluster containing the YSO (Blaauw, 1964; de Geus et al., 1989; Mamajek et al., 2002; Müller et al., 2011; Bossini et al., 2019; Esposito et al., 2020). These objects additionally represent a wide range of .
@@
These data, along with the HST/STIS E140M spectrum of Betelgeuse from Carpenter et al. (2018), are shown in Figure 2. At the time of the Carpenter et al. (2018) observations (conducted in 2011), was transiting Betelgeuse, nearly at the middle of transit (see also § 3.4). Thus, while the overall flux of each emission line would contain contributions from both Betelgeuse and , their relative radial velocities should be close to zero — the only RV shift present at that time should be the overall peculiar motion of the system (22 km/s) as well as any RV changes induced by Betelgeuse’s pulsational or convective activity. In Figure 2, these data are shown as a thick solid orange line (labeled Ori 2011), and have been blueshifted by km/s to match the expected total RV of Betelgeuse during our 2024 observations.
We likewise shift the ULLYSES YSO spectra to be at the expected radial velocity of during the 2024 observations, 45.5 km/s redshifted. We also correct for the distance of Betelgeuse (168 pc, Joyce et al. 2020). To account for extinction, we de-extinct the YSO spectra given their values from the ULLYSES project, and re-extinct them with full extinction curves given the Betelgeuse’s measured extinction. Following Montargès et al. (2021) we adopt values of , appropriate for supergiants (Massey et al., 2005), and 444The of Betelgeuse has been largely consistent across observational works. Early work by Carpenter et al. (1994) fit the GHRS-FUV spectrum with =0.75 and =3. Levesque & Massey (2020) find =0.62 in their 2004-2020 observations. . This is likely a conservative estimate, as any denser gas and dust interior to the orbit of , such as the inner shell at inferred from ALMA observations (e.g. Perrin et al., 2007; Haubois et al., 2019, 2023), would contribute to the extinction of Betelgeuse but not . Extinction parameters are applied to the YSO spectra for the full Milky Way extinction curve (Gordon et al., 2023) using the dust_extinction555https://dust-extinction.readthedocs.io/en/stable/ package (Gordon, 2024).
@@
While the overall flux in the UV of these YSOs is only 1–5% of that of Betelgeuse, several features could be visible in a STIS exposure containing (as opposed to exposures containing only the chromosphere of Betelgeuse). There are 8 strong emission lines in the ULLYSES YSO spectra that are close to or as strong as the same line from Betelgeuse, and 4 that do not appear as emission lines in the spectrum of Betelgeuse. The four panels of Figure 2 highlight four wavelength ranges with promising features that may be visible in our observations of the system.666See Figure 7 and Table 2 of Linsky (2017) for the full suite of lines we examined here and in the observing proposal.
@@
We also note that the ULLYSES data are all of isolated YSO systems. YSOs are often surrounded by, and accrete from, circumstellar disks, leading to the emission “wings” seen in some of the ULLYSES spectra. These are protoplanetary disks at early (10 Myr) times, and gas-poor processed disks at later (10 Myr) times (Ribas et al., 2015). These disks eventually photo-evaporate from high-energy radiation from the central star (Gorti & Hollenbach, 2009). However, these circumstellar disks can also be photoevaporated by UV radiation from nearby OB stars (Adams et al., 2004; Andrews, 2020), which has been directly observed for YSOs in young OB associations (Johnstone et al., 1998). For binary YSOs, systems with circumprimary but no circumsecondary disks are also much more common than the inverse (White & Ghez, 2001), and Kraus et al. (2012) finds that 2/3rds of YSO systems with close separations (r40 AU, compared to the 9AU separation for ) disperse their disks with 1 Myr of formation. Given that the Ori– system has a small separation relative to 40AU, and given that Betelguse was an extremely hot O star during its main-sequence lifetime, there is a high likelihood that a circumstellar disk around the companion has been dispersed by present day. Thus, the wide emission features stemming from strong winds or accretion (red/blue-shifted relative to the companion) that are common in YSO spectra (seen in the ULLYSES spectra in panels (a) and (d) of Figure 2), as well as wings around Lyman-, would almost certainly not be present in the Ori system.
@@
We therefore focus our efforts on narrow lines without those wings, such as fluorescent H2 lines (seen in all four panels of Figure 2). Of the spectra shown here, the 2 M⊙ YSO would be plainly visible in these lines, while the 0.77 M⊙ YSO would never be visible even with a large RV offset between Betelgeuse and . The 0.91 M⊙ YSO also appears generally too faint to detect. The 1.1 and 1.5 M⊙ YSOs are more promising to identify, though the extra flux added to the system is small. We also note the important caveat that these are individual, empirical YSO observations, not theoretical models of YSO FUV chromospheric features across the full range of possible companion masses, for which first-principles calculations remain elusive.
@@ 3 Observations with HST
Below, we summarize the observational history of Betelgeuse to place our campaign in context. We then describe the parameters of our own observational campaign that took place in November 2024. We emphasize the importance of the timing of our observations relative to the companion’s orbital phase (receding quadrature). We also describe our unique aperture tiling pattern, which allows for co-incident spectra which both do and do not contain emission from the companion.
@@ 3.1 Past UV Observations
Long before the launch of HST, early measures of the ultraviolet spectrum of Betelgeuse (spectral type M1-M2) were obtained from a balloon-borne telescope-spectrometer, the OAO–2 spacecraft (Wisconsin photometer-spectrometer), and OAO–3 (Princeton experiment on Copernicus). All detected the Mg II doublet near 2800 Å, a strong emission doublet in the near-ultraviolet (Kondo et al., 1972; Doherty, 1972; Bernat & Lambert, 1976). Subsequently, the International Ultraviolet Explorer (IUE) extended spectroscopy further into the ultraviolet to the Lyman- line near 1200 Å. These low resolution spectra show no evidence of any species of ionization beyond Si II and Fe II, suggesting electron temperatures less than about 10,000 K (Basri et al., 1981). The Hopkins Ultraviolet Telescope on the Space Shuttle offered the possibility of spectroscopy from 800 Å to 1800 Å (Davidsen, 1993); however, the few spectra of Betelgeuse were too low signal-to-noise to be conclusive.
@@
A substantial advance in the study of the near ultraviolet spectrum came with the Goddard High Resolution Spectrograph (GHRS) on HST, acquiring a detailed spectrum from 1980–3000 Å that enabled the identification of many features (Brandt et al., 1995) from single and doubly ionized species, among them Fe II, C II, Ni II, Ti II in addition to molecular emission of OH. The GHRS instrument also accessed the far ultraviolet spectrum, from 1200 Å to 1930 Å, a region that is dominated by CO bandhead absorption in addition to the singly and doubly ionized species, C I, S I, O I, Cr II, and Si II (Carpenter et al., 1994). Those authors also claim the presence of a true continuum, with a color temperature of 3500–4800 K which they ascribe to the chromosphere, and suggest an origin in hot chromospheric regions. A large hot region was later observed on the extended ultraviolet disk in the 2500 Å continuum, with temperatures in excess of 200 K from the disk value (Gilliland & Dupree, 1996), thus supporting origin in a hot region. Carpenter et al. (1994) in fact consider but dismiss the idea that a companion creates the continuum, in part because the observed flux distribution of warmer stars does not match the continuum shape.
@@
The spectral range below 1200 Å, became generally accessible with the launch of the Far Ultraviolet Spectroscopic Explorer (FUSE), which enabled access from 905–1179 Å where highly ionized species are expected. However, a spectrum of Betelgeuse revealed no sign of C III or O VI which are found in giants and supergiants of spectral type F, G, and K (Dupree et al., 2005).
@@
More recently, observations of Betelgeuse (PI T. Ayres 2010) were performed with STIS E140M using a square arcsec aperture for 4.5 hours (Carpenter et al., 2018).The dates of these observations were February 7, 8, 24, 26, and April 1st 2011, at which point was transiting (passing in front of) Betelgeuse. While our campaign used the same instrument, grating, and aperture as the 2011 campaign, ours differs critically in timing, tiling, and motivation. The previous Ayres (2010); Carpenter et al. (2018) campaign was only interested in the overall UV spectrum of the chromosphere, and they did not vary their pointing. Therefore, that campaign did not obtain any spatial information and did not collect any “clean” spectra without .